OpenStudy (empty):

Let's learn and try to prove the ABC conjecture! (No prior knowledge of Inter-Universial Teichmuller theory necessary!)

OpenStudy (empty):

Ok what the hell is the ABC conjecture? First we have to define a function, \[rad(n)\] All it does is gives you a number with the exponents on its prime factorization thrown away leaving only 1s left, like this: \[rad(12)=rad(2^23^1)=2^13^1=6\] Now we're almost ready to state the conjecture. We start with two numbers \(a\) and \(b\), \[a+b=c\] and here's the conjecture! \[c> rad(abc)^k\] This statement is true for only finitely many numbers for any choice of \(k>1\)

OpenStudy (empty):

@FireKat97 Ok I'm not sure I fully understand this I guess I'm going to try throwing in random examples to play with it to try to figure it out. :P

OpenStudy (empty):

Ok something interesting is if a and b don't share factors, then neither does c, so they're all relatively prime: \[\gcd(a,b)=\gcd(a,c)=\gcd(b,c)=1\] Since that's true I think we can separate out the rad function this way since it won't matter: \[rad(abc)=rad(a)*rad(b)*rad(c)\]

OpenStudy (empty):

Anybody interested in trying to figure this out or have any ideas of interesting stuff? I don't wanna feel like I'm yelling in an echo chamber lol.

OpenStudy (firekat97):

does it help if I say 'everything above makes sense'? xD

OpenStudy (dan815):

what do u mean its only true for finitely many numbers for some choice k > 1

OpenStudy (dan815):

arent there an infinite nuumber for c, that we can pick

OpenStudy (kainui):

Prove it dan, then you'll have found a counter example to the abc conjecture lol

OpenStudy (dan815):

like suppose i pick c^n+1, and a=1,b=1, then rad(abc)= c and c^n+1 = c^k now i can pick n=k and get an infinite number right

OpenStudy (kainui):

No because a+b=c

OpenStudy (dan815):

ohhhhhhhhhhh

OpenStudy (dan815):

ok now i see why there could be finite okay!

OpenStudy (kainui):

Haha yeah I feel like this is related to the arithmetic derivative, also did you notice the rad(n) function is multiplicative? Another fun fact although probably useless is this identity you could call it with the mobius function. \[\mu(rad(n)) \ne 0\]

OpenStudy (dan815):

true

OpenStudy (dan815):

lol

OpenStudy (kainui):

I was looking at this \[c=p^{k+1}\] so that we have \[p > rad^k(ab)\]

OpenStudy (dan815):

oo ok ok

OpenStudy (dan815):

a + b = c and a*b <c

OpenStudy (kainui):

I guess this probably always false maybe I should have just written this out: \[c=p^k\] \[p^k > rad^k(abp^k)=rad^k(ab) p^k\] \[1 > rad(ab)\] So like because of stuff like this there are infinitely many cases where I see that it must be false it's just proving that there are finitely many cases where it's true, kinda weird.

OpenStudy (dan815):

c=p^(k+1) p > (ab)^k

OpenStudy (kainui):

That's only true if a and b are primes

OpenStudy (kainui):

\[rad^y(p^x)=p^y\] It's kinda like a prime renamer huh interesting.

OpenStudy (dan815):

a+b = p^(k+1) p=root_{k+1} (a+b) <--- this being an integer has some big restrictions

OpenStudy (kainui):

wait what do you mean by root_{k+1}

OpenStudy (dan815):

K+1 th root of A+B

OpenStudy (kainui):

``` \[ \sqrt[k+1]{a+b}\] ``` \[ \sqrt[k+1]{a+b}\]

OpenStudy (dan815):

yea lol

OpenStudy (kainui):

Oh it has some big restrictions but it's not a big deal that it does.

OpenStudy (kainui):

Just start here: \[p^{k}=\frac{p^k+r}{2} + \frac{p^k-r}{2}\] there are always going to be multiple choices of r that work.

OpenStudy (dan815):

u there?

OpenStudy (kainui):

Yeah I'm here I was watching some youtubes what's up

OpenStudy (dan815):

why are you doing that expression with r

OpenStudy (kainui):

I think I kinda made that weirder than it needed to be, I was just showing we can always find a value of a and b that work, \[c=p^k\]\[a=p^k+r\]\[b=p^k-r\] as long as r isn't divisible by p, then all 3 numbers are relatively prime.

OpenStudy (dan815):

ohh i see! ok ok

OpenStudy (kainui):

cause you were trying to say that there are restrictions or whatever. I guess I was looking for an infinite number of counter examples I think I'm partly confused too haha

OpenStudy (dan815):

/2

OpenStudy (kainui):

Yeah ok the divide by 2 has to be in there right I forgot

OpenStudy (dan815):

okay hmm and this is can even be the maximum one to find

OpenStudy (dan815):

this should produce the largest or close to largest answers from

OpenStudy (dan815):

rad(abc) for small rs

OpenStudy (kainui):

I was thinking that but I realized it's flawed

OpenStudy (kainui):

cause it's not so much the size of the number it's the amount of unique primes in the number that matters more

OpenStudy (dan815):

oh thats ttrue

OpenStudy (kainui):

Also this way isn't really the only way we can pick numbers a and b where they're some distance from the middle, we could pick like \[p^k = (p^k-q^r)+(q^r)\] or something like this idk

OpenStudy (kainui):

Ooooh ok I got an idea \[p^k = q \# + (p^k-q \#)\] where q# is the largest primorial not exceeding \(p^k\) this should help us maximize the rad function!

OpenStudy (dan815):